OpenStudy (iwanttogotostanford):

SIMPLIFY HELP< KEEP GETTING STUCK: Simplify: (2x-5)^2

OpenStudy (iwanttogotostanford):

don't need answer but explanation

OpenStudy (iwanttogotostanford):

I actually need help explaining it like where'd you get the -20x

OpenStudy (anonymous):

I want to be a math calculator so bad i just need more medals

OpenStudy (anonymous):

k(2x−5)2 =(2x+−5)(2x+−5) =(2x)(2x)+(2x)(−5)+(−5)(2x)+(−5)(−5) =4x2−10x−10x+25 =4x2−20x+25

OpenStudy (anonymous):

does that help?

OpenStudy (ryuga):

@pooja195

OpenStudy (iwanttogotostanford):

oh ok so I use F.O.I.L?

OpenStudy (anonymous):

pooja is busy with a drama chat

OpenStudy (anonymous):

yup foil

OpenStudy (iwanttogotostanford):

oh ok thanks

OpenStudy (anonymous):

next time you got a question LINK ME IN IT k id love to help

OpenStudy (iwanttogotostanford):

what is 5x^3 plus x^3 ? @Lord_Frieza

OpenStudy (anonymous):

Ive been given a warning for helping you out

OpenStudy (iwanttogotostanford):

ya because you just gave me the answers...

OpenStudy (anonymous):

k let me explain the question and the steps